OpenStudy (anonymous):

Suppose you deposit $5,000 in a savings account that earns 3% annual interest. If you make no other withdrawals or deposits, how many years will it take the account balance to reach at least $6,000?

OpenStudy (luigi0210):

Use: \[\LARGE A= P(1 + r)^n \] So: \[\LARGE 6000=5000(1+.03)^n \] n=number of years a=amount p=principal amount r=rate

OpenStudy (anonymous):

I got n = 7? 6149.369327

OpenStudy (luigi0210):

Yup, so approx. 7 years.